Bella is wrapping a cube-shaped gift box that measures 8.5 inches along each edge. How many square inches of wrapping paper is needed to cover the box completely without overlapping? Show work
Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20056)

The cube has 6 square 8.5×8.5 sides, or faces:

1. A front side, or 8.5×8.5=72.5 square inch front side face
2. A back side, or 8.5×8.5=72.5 square inch back side face
3. A left side, or 8.5×8.5=72.5 square inch left side face
4. A right side, 8.5×8.5=72.5 square inch or right side face
5. A top side, or 8.5×8.5=72.5 square inch top side face
6. A bottom side, or  8.5×8.5=72.5 square inch side face face


So we multiply 72.5 square inches by 6 and get that 433.5 
square inches of wrapping paper would cover the box completely 
without overlapping.

Edwin

Answer by ikleyn(52788)

.
SOLUTION.  6*8.5*8.5 = 6*72.25 = 433.5 square inches.    ANSWER


It is not always the longer solution is more understandable, more correct and/or more justified.

In opposite, the longer solution may contain more mistakes.